OpenStudy (anonymous):

2p+5q-r=25 2p-2q-r=-24 3p-q+5r=4

Mehek (mehek14):

are those separate questions? and are you solving for p and q? @Amandasteele

OpenStudy (anonymous):

It's all one question and yes solving for both p and q

OpenStudy (gorv):

take first 2 equation

OpenStudy (gorv):

we need p and q

OpenStudy (gorv):

so subtract first 2 from each other and tell me what u got

OpenStudy (gorv):

tell me what u got ??

OpenStudy (gorv):

u will get value of q from here

OpenStudy (gorv):

subtract and post what u got ??

OpenStudy (anonymous):

To solve this kind of simultaneous equation, you make one of the variables of the first equation as the subject of the formula. Once you find it, replace such variable in the remaining 2 equations.Thirdly, you make one of the 2 remaining variables as the subject of the formula to replace it in the last equation and surely you will find the value of one of the variables. Fourthly, you will move backwards, by going to the previous equation to replace the found variable to find the other one and repeat this with the next previous equation and you shall get the value as well. Mathematically I am saying:

OpenStudy (anonymous):

\[25p+5p-25=r\]

OpenStudy (anonymous):

Replacing it in the 2 other equations...\[2p-2q-2p-5q+25+24=0\]

OpenStudy (anonymous):

and \[3p-q+5(2p+5q-25)-4=0\]

OpenStudy (unklerhaukus):

can we use a matrix?

OpenStudy (anonymous):

If you work out the equation \[2p-2q-2p-5q+25+24=0\], q=7, which you will replace it in the last equation, of \[3p-7+5(2p+5*7-25)-4=0\], which simplifies to \[-3p-7+10p+50-4=0\] and \[p=-\frac{ 39 }{ 7 }\].

OpenStudy (anonymous):

Now that we have P and Q, let us replace it in the very first equation where we made r the subject. \[2(-\frac{ 39 }{ 7 })+5(7)-25=r\] and r=-8/7

OpenStudy (anonymous):

There. All the variables were found.

OpenStudy (anonymous):

@UnkleRhaukus , I am not familiar with the matrix method that finds 3 variables of polynomials.

OpenStudy (unklerhaukus):

\[2p+5q-r=\ \ \ 25\\ 2p-2q-r=-24\\ 3p-q+5r=\ \ \ \ \ 4\] \[\begin{bmatrix}2&5&-1\\2&-2&-1\\3&-1&5\end{bmatrix} \begin{bmatrix}p\\q\\r\end{bmatrix} =\begin{bmatrix}\ \ \ 25\\-24\\\ \ \ \ \ 4\end{bmatrix}\] \[\left[\begin{array}{rrr|r}2&5&-1&25\\2&-2&-1&-24\\3&-1&5&4\end{array}\right]\]

OpenStudy (unklerhaukus):

\[R_1\to R_3-R_1\]\[\left[\begin{array}{rrr|r}1&-6&6&-21\\2&-2&-1&-24\\3&-1&5&4\end{array}\right]\] \[R_2\to R_2-2R_1\] \[\left[\begin{array}{rrr|r}1&-6&6&-21\\0&10&-13&18\\3&-1&5&4\end{array}\right]\]

OpenStudy (anonymous):

I think I recall a bit. It looks like Gaussian method.

OpenStudy (unklerhaukus):

\[R_2\to R_2/10\]\[R_3\to R_3-3R_1\]\[\left[\begin{array}{rrr|r}1&-6&6&-21\\0&1&-1.3&1.8\\0&17&-13&67\end{array}\right]\]

OpenStudy (unklerhaukus):

\[R_1\to R_1+6R_2\]\[R_3\to R_3-17R_2\]\[\left[\begin{array}{rrr|r}1&0&-1.8&10.2\\0&1&-1.3&1.8\\0&0&9.1&36.4\end{array}\right]\]

OpenStudy (unklerhaukus):

\[R_3\to R_3/9.1\]\[\left[\begin{array}{rrr|r}1&0&-1.8&10.2\\0&1&-1.3&1.8\\0&0&1&4\end{array}\right]\] \[R_1\to R_1+1.8R_3\\ R_2\to R_2+1.3 R_3\]\[\left[\begin{array}{rrr|r}1&0&0&17.4\\0&1&0&7\\0&0&1&4\end{array}\right]\] \[\begin{bmatrix}p\\q\\r\end{bmatrix} =\begin{bmatrix}17.4\\ 7 \\4\end{bmatrix}\] hmm `17.4`, i probably made an error, lets check the result anyway...

OpenStudy (anonymous):

What we had in common was q.

OpenStudy (unklerhaukus):

\[2p+5q-r=\ \ \ 25\\ 2p-2q-r=-24\\ 3p-q+5r=\ \ \ \ \ 4\\~\\\qquad\qquad\downarrow\\~\\~ 2(17.4)+5(7)-(4)=65.8\\ 2(17.4)-2(7)-(4)=16.8\\ 3(17.4)-(7)+5(4)=65.2\] errors \(\ddot \frown\)

OpenStudy (unklerhaukus):

i bet p does not equal 17.4 \[3p-(7)-5(4)=4\] \[3p=4+7-20=-9\\ p=-3\]

OpenStudy (anonymous):

Oh my ... We have to retry this. I am restarting now...

OpenStudy (unklerhaukus):

do these values solve the system? \[2p+5q-r=\ \ \ 25\\ 2p-2q-r=-24\\ 3p-q+5r=\ \ \ \ \ 4\\~\\\qquad\qquad\downarrow\\~\\~ 2(-3)+5(7)-(4)=25\\ 2(-3)-2(7)-(4)=-24\\ 3(-3)-(7)+5(4)=4\]

OpenStudy (unklerhaukus):

B I N G O !

OpenStudy (anonymous):

Right, good!

OpenStudy (unklerhaukus):

did you get your method to work @Hoslos ?

OpenStudy (anonymous):

Haha. You know? I re-did it and found new values. when plotted in the first 2 equations it went correct, but in the third one did not. Strange. it is my first time missing this method... Maybe because it is a 3-term variable expression, whereas I used this method a lot with 4-term expressions.

OpenStudy (unklerhaukus):

i found my error ... \[R_1\to R_1+6R_2\\\left[\begin{array}{rrr|r}1&0&-1.8&\color{red} -10.2\\0&1&-1.3&1.8\\0&0&9.1&36.4\end{array}\right]\] ... \[R_1\to R_1+1.8R_3\\ \left[\begin{array}{rrr|r}1&0&0&-3\\0&1&0&7\\0&0&1&4\end{array}\right]\]